# Wellington Road, Clayton, Victoria 3800, Australia

### v1

@@ 1 Introduction
While considerable attention has been devoted to the theories of ADM, its detectability in particle colliders remains relatively less well-explored. A previous study by March-Russell et al. (2012) March-Russell:2012elz presented collider and DD constraints on ADM, together with the constraints imposed by the model itself. The key idea being that efficient annihilation of the symmetric part of the ADM relic density requires that the coupling between DM and SM fermions cannot be too weak. This is under the usual simplifying assumption that the dark matter annihilates into and scatters off SM fermions, and these processes are mediated by the same elementary interaction. Quite remarkably, already in this early work, it was found that collider data and direct detection searches practically rule out ADM with mass all the way up to 100 GeV or more.
@@
We then turn our attention to a leptophilic variant of the ADM scenario, where ADM couples to SM leptons but not to quarks. Such ADM models may be motivated by theory, and are poised to take advantage of the upcoming experiments. In some ADM models, the DM asymmetry fundamentally arises due to new sphalerons that transform the lepton asymmetry created by leptogenesis into a DM asymmetry. Here, a chiral gauge interaction shared by DM and leptons directly leads to the above phenomenology and does not necessarily need to involve quarks. On the observational side, DM interactions with leptons (electrons, in particular) has gained interest over the past decade. Unsurprisingly, DM-lepton interactions have been studied in depth  Baltz:2002we ; Chen:2008dh ; Fox:2008kb ; Cao:2009yy ; Cohen:2009fz ; Ibarra:2009bm ; Bi:2009uj ; Kopp:2009et ; Ko:2010at ; Chao:2010mp ; Schmidt:2012yg ; Das:2013jca ; Schwaller:2013hqa ; Bai:2014osa ; Agrawal:2014ufa ; Kopp:2014tsa and a gamut of experiments have already produced interesting limits on leptophilic DM XENON100:2015tol ; LZ:2021xov ; Chen:2018vkr ; Marsicano:2018vin . More importantly, there are now several novel proposals that promise to make significant progress in this direction in the near future CRESST:2017ues ; SuperCDMS:2020aus ; SENSEI:2020dpa ; DAMIC-M:2023gxo ; aguilar2022oscura ; Marshall:2020azl ; Hertel:2018aal ; SPICE:2023aqd ; 2020JLTP ; Hochberg:2015pha ; Hochberg:2019cyy ; Geilhufe:2019ndy ; Kahn:2021ttr ; Blanco:2022cel ; Huang:2023nms ; Das:2023cbv ; Carew:2023qrj . Limits on DM-lepton interactions from the heating of neutron stars (NS) Bell:2019pyc and white dwarfs (WD) Bell:2021fye happen to provide some of the strongest constraints. On the collider front, mono-photon searches at LEP have been set bounds on four effective operators (three of which are independent) Fox:2011fx , which are more stringent than the NS or WD bounds for DM masses in the 1-100 GeV range. Investigations have also focused on the prospects of identifying leptophilic dark matter through future lepton colliders, mainly the International Linear Collider (ILC) Dutta:2017ljq ; Barman:2021hhg ; Kundu:2021cmo . In the latter part of this paper, we update the bounds on DM-lepton interactions using the LEP mono-photon searches, compare with other constraints as for the quark case, and appraise the viability of leptophilic ADM. We also forecast discovery prospects at the prospective FCC-ee.
@@ 2.1 Effective interactions
We assume the DM is a Dirac fermion and the mediators responsible for its interactions with the SM particles are heavier than the DM itself. The interactions are thus encoded in higher-dimensional operators within the framework of an effective field theory (EFT) involving SM and DM degrees of freedom. However, we will ignore the dimension-5 couplings of DM to the SM Higgs, which is already very well studied and constrained. See, for example, the stringent upper limits of $\rm H_{SM}\to invisible$ branching fractions CMS:2022qva ; ATLAS:2022yvh . [^1] Note that the Higgs portal for fermionic DM of mass smaller than $\sim$20 GeV, consistent with $\rm BR(H_{SM}\to invisible)<10\%$, leads to DM-nucleon scattering cross-section below the “neutrino floor” Strigari:2009bq ; Billard:2013qya ; Monroe:2007xp , which is uninteresting from DD perspective Arcadi:2021mag . Alternatively, though a higher mass region of DM allows parameter regions above the neutrino floor, they are excluded by the DD limits Arcadi:2021mag . With these considerations, in Table 1 we write down all allowed EFT operators at dimension-6. These operators are of the form
@@
Among the tensor operators, $\bar{\psi}\sigma^{\mu\nu}\psi\bar{f}i\sigma_{\mu\nu}\gamma^{5}f$ is rewritable as $\bar{\psi}i\sigma^{\mu\nu}\gamma^{5}\psi\bar{f}\sigma_{\mu\nu}f\equiv\mathcal{O}_{pt}$ and $\bar{\psi}i\sigma^{\mu\nu}\gamma^{5}\psi\bar{f}i\sigma_{\mu\nu}\gamma^{5}f$ as $\bar{\psi}\sigma^{\mu\nu}\psi\bar{f}\sigma_{\mu\nu}f\equiv\mathcal{O}_{tt}$. So they are not separately considered. Moreover, operators enter the Lagrangian along with respective Wilson coefficients (WC). In this work, we will consider that the WCs are absorbed into the $\Lambda$.
@@ 2.2 Relic density and the asymmetry criterion
The asymmetric dark matter paradigm assumes that both DM ($\chi$) and the anti-DM ($\bar{\chi}$) are present up to some era of the evolution of the Universe, as in the case of baryonic matter. Then by some mechanism, an asymmetry is generated between DM and anti-DM sectors, which gets frozen out with the departure from thermal equilibrium, and the symmetric part gets annihilated out. In this study we remain agnostic about the asymmetry generation and sharing mechanisms. The total dark-abundance can be written as the sum of the yields of DM and anti-DM, i.e., $Y_{\chi}+Y_{\overline{\chi}}$, where the yield $Y=n/s$ is defined as the number density scaled by the entropy density of the early universe. This, in turn can be expressed as a symmetric and an asymmetric component:
@@
with $x_{\rm F_{0}}=m_{\chi}/T_{\rm F}$ being the usual coordinate at freeze-out. The difference between $x_{\rm F}$ and $x_{\rm F_{0}}$ encodes the ADM correction to the $\bar{\chi}$ abundance, which is otherwise under-predicted by the standard WIMP treatment. The expressions of the annihilation cross-sections times velocity for each of the operators in Table 1 are given in Appendix A. The constraint imposed in ADM model is that the symmetric part, as in Eq. 4, must be decidedly sub-leading to the asymmetric part. We implement this by restricting $Y_{\rm sym}$ to contribute $\leq 1\%$ (say) to $\Omega_{\rm DM}h^{2}$ March-Russell:2012elz :
@@ 3 Constraining ADM-Quark Interactions
In this section, we first revisit the constraints on the DM-quark effective interactions and then find their implications on ADM scenario. The DM-quark interactions are mainly constrained by the DD experiments and the results of mono-jet searches at the LHC, where direct interaction of DM with quarks is present. These constraints were already shown in Ref. March-Russell:2012elz . More extensive experimental results and several new searches have come up since, which are expected to impose stronger and more robust constraints. In this section, we update the bounds derived from mono-jet searches at LHC using latest data and by taking relevant detector effects into account. We then present the big picture for ADM-quark interactions, comparing with constraints from the asymmetry, and including other important constraints, especially those from recent DD experiments.
@@ 4.1 Mono-photon searches at LEP
Now we use this setup to generate the mono-photon energy distribution for the mono-photon production for DM (Eq. 12) for each choice of $m_{\chi}$ and $\Lambda$. Calculation of $\chi^{2}$ between the photon energy distributions for SM-MC and the DM, it is possible to predict the allowed and ruled out DM parameter space. Results are shown for different operators in Figs. 10-14. Notice that, these bounds are the strongest one for leptophilic dark matter in the 1-100 GeV mass range, which are discussed later.
@@ 4.4 Constraints and discovery potential for leptophilic ADM
We present the overall picture of the allowed and excluded parameter spaces for DM-electron interaction in Figs. 10 -14. Each of the shaded regions within the presented figures signify exclusion due to specific hypotheses. For instance, the blue shaded region eliminates the possibility of dark matter being asymmetric, while the dark red portion is excluded based on the outcomes of the LEP mono-photon searches, and so on. Interestingly, for $\mathcal{O}_{ss},\mathcal{O}_{sp},\mathcal{O}_{aa},\mathcal{O}_{av}$, where $\chi\bar{\chi}\to e^{+}e^{-}$ occurs through velocity suppressed $p$-wave ($s$-wave annihilation is either absent or has $m_{q}^{2}$ suppression), the ADM requirement excludes a larger parameter space in the $\Lambda-m_{\chi}$ plane. It is important to note here that the other exclusions, which are not restricted to the ADM hypothesis, hold for any DM with these effective interactions, and all the regions above the shaded region (except blue) are allowed. Our study reveals that, though the constraints from the NS or WDs on the DM-lepton interactions are overall strong, the bounds coming from the mono-photon searches at the LEP experiments surpasses all those in the sensitive region, $m_{\chi}\sim 1-100$ GeV. Considering the ADM hypothesis, we find that most of the operators fail to provide an ADM solution for any value of $\Lambda$ in the region $m_{\chi}\sim 1-100$ GeV, sensitive to the LEP experiment. Only for operators $\mathcal{O}_{vv}$, $\mathcal{O}_{av}$, and $\mathcal{O}_{ps}$, the bounds are somewhat weak $\sim$ 50 GeV, while for $\mathcal{O}_{va}$, $\mathcal{O}_{pt}$, $\mathcal{O}_{sp}$, and $\mathcal{O}_{pp}$ they are much weaker $\sim$10 GeV. Notice that, for the interactions $\mathcal{O}_{ss}$, $\mathcal{O}_{sp}$, $\mathcal{O}_{ps}$, and $\mathcal{O}_{pp}$ the LEP bounds are the only ones that can completely exclude ADM possibility of a certain mass.
@@ 5 Summary and Outlook
Next, we concentrate on the ADM-lepton interactions. In this case, the strongest bounds generally come from NS or WD studies. However, we found that the constraints from the monophoton searches at the LEP experiment, which we have calculated in detail for each interaction, overshoot those from the NS or WD for all types of interactions in the 1-100 GeV mass range. According to our study, the possibility of a leptophilic ADM is ruled out by the LEP experiment for $\sim$ 1-100 GeV mass range of the ADM, except for the vector--axial-vector operator ($\mathcal{O}_{va}\equiv$ $\frac{1}{\Lambda^{2}}\bar{\psi}\gamma^{\mu}\psi\bar{e}\gamma_{\mu}\gamma^{5}e$), the pseudo-tensor--tensor operator ($\mathcal{O}_{pt}\equiv$$\frac{1}{\Lambda^{2}}\bar{\psi}\sigma^{\mu\nu}i\gamma^{5}\psi\bar{e}\sigma_{\mu\nu}e$), the scalar--pseudo-scalar operator ($\mathcal{O}_{sp}\equiv\frac{1}{\Lambda^{2}}\bar{\psi}\psi\bar{e}i\gamma^{5}e$), and the pseudo-scalar--pseudo-scalar operator ($\mathcal{O}_{pp}\equiv\frac{1}{\Lambda^{2}}\bar{\psi}\gamma^{5}\psi\bar{e}\gamma^{5}e$) mediated interactions, where the constraints are loose $\sim$ $\mathcal{O}$(10) GeV. Again, the constraints apply to any DM having the considered effective interactions (Table 1).

### v2

@@ 1 Introduction
While considerable attention has been devoted to the theories of ADM, its detectability in particle colliders remains relatively less well-explored. Two previous studies, by Buckley Buckley:2011kk and by March-Russell et al. (2012) March-Russell:2012elz , presented collider and DD constraints on ADM, together with the constraints imposed by the model itself. The key idea being that efficient annihilation of the symmetric part of the ADM relic density requires that the coupling between DM and SM fermions cannot be too weak. This is under the usual simplifying assumption that the dark matter annihilates into and scatters off SM fermions, and these processes are mediated by the same elementary interaction. Quite remarkably, already in this early work, it was found that collider data and direct detection searches practically rule out ADM with mass all the way up to 100 GeV or more.
@@
We then turn our attention to a leptophilic variant of the ADM scenario, where ADM couples to SM leptons but not to quarks. Such ADM models may be motivated by theory, and are poised to take advantage of the upcoming experiments. In some ADM models, the DM asymmetry fundamentally arises due to new sphalerons that transform the lepton asymmetry created by leptogenesis into a DM asymmetry. Here, a chiral gauge interaction shared by DM and leptons directly leads to the above phenomenology and does not necessarily need to involve quarks. On the observational side, DM interactions with leptons (electrons, in particular) has gained interest over the past decade. Unsurprisingly, DM-lepton interactions have been studied in depth  Baltz:2002we ; Chen:2008dh ; Fox:2008kb ; Cao:2009yy ; Cohen:2009fz ; Ibarra:2009bm ; Bi:2009uj ; Kopp:2009et ; Ko:2010at ; Chao:2010mp ; Schmidt:2012yg ; Das:2013jca ; Schwaller:2013hqa ; Bai:2014osa ; Agrawal:2014ufa ; Kopp:2014tsa and a gamut of experiments have already produced interesting limits on leptophilic DM XENON100:2015tol ; LZ:2021xov ; Chen:2018vkr ; Marsicano:2018vin . More importantly, there are now several novel proposals that promise to make significant progress in this direction in the near future CRESST:2017ues ; SuperCDMS:2020aus ; SENSEI:2020dpa ; DAMIC-M:2023gxo ; aguilar2022oscura ; Marshall:2020azl ; Hertel:2018aal ; SPICE:2023aqd ; 2020JLTP ; Hochberg:2015pha ; Hochberg:2019cyy ; Geilhufe:2019ndy ; Kahn:2021ttr ; Blanco:2022cel ; Huang:2023nms ; Das:2023cbv ; Carew:2023qrj . Limits on DM-lepton interactions from the heating of neutron stars (NS) Bell:2019pyc and white dwarfs (WD) Bell:2021fye happen to provide some of the strongest constraints. On the collider front, mono-photon searches at LEP have been set bounds on four effective operators (three of which are independent) Fox:2011fx , which are more stringent than the NS or WD bounds for DM masses in the $\lesssim 100$ GeV range. Investigations have also focused on the prospects of identifying leptophilic dark matter through future lepton colliders, mainly the International Linear Collider (ILC) Dutta:2017ljq ; Barman:2021hhg ; Kundu:2021cmo . In the latter part of this paper, we update the bounds on DM-lepton interactions using the LEP mono-photon searches, compare with other constraints as for the quark case, and appraise the viability of leptophilic ADM. We also forecast discovery prospects at the prospective FCC-ee.
@@ 2.1 Effective interactions
We assume the DM is a Dirac fermion and the mediators responsible for its interactions with the SM particles are heavier than the DM itself. The interactions are thus encoded in higher-dimensional operators within the framework of an effective field theory (EFT) involving SM and DM degrees of freedom Beltran:2010ww ; Goodman:2010ku . However, we will ignore the dimension-5 couplings of DM to the SM Higgs, which is already very well studied and constrained. See, for example, the stringent upper limits of $\rm H_{SM}\to invisible$ branching fractions CMS:2022qva ; ATLAS:2022yvh . [^1] Note that the Higgs portal for fermionic DM of mass smaller than $\sim$20 GeV, consistent with $\rm BR(H_{SM}\to invisible)<10\%$, leads to DM-nucleon scattering cross-section below the “neutrino floor” Strigari:2009bq ; Billard:2013qya ; Monroe:2007xp , which is uninteresting from DD perspective Arcadi:2021mag . Alternatively, though a higher mass region of DM allows parameter regions above the neutrino floor, they are excluded by the DD limits Arcadi:2021mag . With these considerations, in Table 1 we write down all allowed EFT operators at dimension-6. These operators are of the form
@@
Among the tensor operators, $\bar{\psi}\sigma^{\mu\nu}\psi\bar{f}i\sigma_{\mu\nu}\gamma^{5}f$ is rewritable as $\bar{\psi}i\sigma^{\mu\nu}\gamma^{5}\psi\bar{f}\sigma_{\mu\nu}f\equiv\mathcal{O}_{pt}$ and $\bar{\psi}i\sigma^{\mu\nu}\gamma^{5}\psi\bar{f}i\sigma_{\mu\nu}\gamma^{5}f$ as $\bar{\psi}\sigma^{\mu\nu}\psi\bar{f}\sigma_{\mu\nu}f\equiv\mathcal{O}_{tt}$. So they are not separately considered. Moreover, operators enter the Lagrangian along with respective Wilson coefficients (WC). In this work, we will consider that the WCs are absorbed into the $\Lambda$. It is to be noted that while $\Lambda$ generally sets the upper limit for momentum transfer where the operator expansion in the EFT remains reliable, it does not establish a strict cut-off. A precise determination of EFT cut-off requires detailed knowledge of the masses and interaction strengths within the UV theory Busoni:2013lha ; Contino:2016jqw . Nevertheless, a rough estimation of EFT validity is always feasible. For instance, assuming maximal coupling in the UV theory, one can derive the condition for validity of the effective interactions. For example, in our case it turns out to be $\Lambda\gtrsim m_{\chi}/2\pi$ with an assumption of s-channel momentum transfer Busoni:2013lha ; Shoemaker:2011vi .
@@ 2.2 Relic density and the asymmetry criterion
The asymmetric dark matter paradigm assumes that both DM ($\chi$) and the anti-DM ($\bar{\chi}$) are present up to some era of the evolution of the Universe, as in the case of baryonic matter. Then by some mechanism, an asymmetry is generated between DM and anti-DM sectors, which gets frozen out with the departure from thermal equilibrium, and the symmetric part gets annihilated out. In this study we remain agnostic about the asymmetry generation and sharing mechanisms.
The total dark-abundance can be written as the sum of the yields of DM and anti-DM, i.e., $Y_{\chi}+Y_{\overline{\chi}}$, where the yield $Y=n/s$ is defined as the number density scaled by the entropy density of the early universe. This, in turn can be expressed as a symmetric and an asymmetric component Graesser:2011wi ; Iminniyaz:2011yp :
@@
with $x_{\rm F_{0}}=m_{\chi}/T_{\rm F}$ being the usual coordinate at freeze-out. The difference between $x_{\rm F}$ and $x_{\rm F_{0}}$ encodes the ADM correction to the $\bar{\chi}$ abundance, which is otherwise under-predicted by the standard WIMP treatment. The expressions of the annihilation cross-sections times velocity for each of the operators in Table 1 are given in Appendix A.
The constraint imposed in ADM model is that the symmetric part, as in Eq. 4, must be decidedly sub-leading to the asymmetric part. We implement this by restricting $Y_{\rm sym}$ to contribute $\leq 1\%$ (say) to $\Omega_{\rm DM}h^{2}$ March-Russell:2012elz :
@@ 3 Constraining ADM-Quark Interactions
In this section, we first revisit the constraints on the DM-quark effective interactions and then find their implications on ADM scenario. The DM-quark interactions are mainly constrained by the DD experiments and the results of mono-jet searches at the LHC, where direct interaction of DM with quarks is present. These constraints were already shown in Ref. Buckley:2011kk ; March-Russell:2012elz . More extensive experimental results and several new searches have come up since, which are expected to impose stronger and more robust constraints. In this section, we update the bounds derived from mono-jet searches at LHC using latest data and by taking relevant detector effects into account. We then present the big picture for ADM-quark interactions, comparing with constraints from the asymmetry, and including other important constraints, especially those from recent DD experiments.
@@ 3.2 Direct detection experiments
where $\Delta_{q}^{p}$ accounts for the spin content of the nucleon and $(\sum_{q}\Delta_{q}^{p})^{2}\simeq 0.32$ COMPASS:2007esq .
@@ 4.1 Mono-photon searches at LEP
Now we use this setup to generate the mono-photon energy distribution for the mono-photon production for DM (Eq. 12) for each choice of $m_{\chi}$ and $\Lambda$. Calculation of $\chi^{2}$ between the photon energy distributions for SM-MC and the DM, it is possible to predict the allowed and ruled out DM parameter space. Results are shown for different operators in Figs. 10-14. Notice that, these bounds are the strongest one for leptophilic dark matter in the $\lesssim 100$ GeV mass range, which are discussed later.
@@ 4.4 Constraints and discovery potential for leptophilic ADM
We present the overall picture of the allowed and excluded parameter spaces for DM-electron interaction in Figs. 10 -14. Each of the shaded regions within the presented figures signify exclusion due to specific hypotheses. For instance, the blue shaded region eliminates the possibility of dark matter being asymmetric, while the dark red portion is excluded based on the outcomes of the LEP mono-photon searches, and so on. Interestingly, for $\mathcal{O}_{ss},\mathcal{O}_{sp},\mathcal{O}_{aa},\mathcal{O}_{av}$, where $\chi\bar{\chi}\to e^{+}e^{-}$ occurs through velocity suppressed $p$-wave ($s$-wave annihilation is either absent or has $m_{q}^{2}$ suppression), the ADM requirement excludes a larger parameter space in the $\Lambda-m_{\chi}$ plane. It is important to note here that the other exclusions, which are not restricted to the ADM hypothesis, hold for any DM with these effective interactions, and all the regions above the shaded region (except blue) are allowed. Our study reveals that, though the constraints from the NS or WDs on the DM-lepton interactions are overall strong, the bounds coming from the mono-photon searches at the LEP experiments surpasses all those in the sensitive region, $m_{\chi}\lesssim 100$ GeV. Considering the ADM hypothesis, we find that most of the operators fail to provide an ADM solution for any value of $\Lambda$ in the region $m_{\chi}\lesssim 100$ GeV, sensitive to the LEP experiment. Only for operators $\mathcal{O}_{vv}$, $\mathcal{O}_{av}$, and $\mathcal{O}_{ps}$, the bounds are somewhat weak $\sim$ 50 GeV, while for $\mathcal{O}_{va}$, $\mathcal{O}_{pt}$, $\mathcal{O}_{sp}$, and $\mathcal{O}_{pp}$ they are much weaker $\sim$10 GeV. Notice that, for the interactions $\mathcal{O}_{ss}$, $\mathcal{O}_{sp}$, $\mathcal{O}_{ps}$, and $\mathcal{O}_{pp}$ the LEP bounds are the only ones that can completely exclude ADM possibility of a certain mass.
@@ 5 Summary and Outlook
Next, we concentrate on the ADM-lepton interactions. In this case, the strongest bounds generally come from NS or WD studies. However, we found that the constraints from the monophoton searches at the LEP experiment, which we have calculated in detail for each interaction, overshoot those from the NS or WD for all types of interactions in the $\lesssim$100 GeV mass range. According to our study, the possibility of a leptophilic ADM is ruled out by the LEP experiment up to 100 GeV mass range of the ADM, except for the vector--axial-vector operator ($\mathcal{O}_{va}\equiv$ $\frac{1}{\Lambda^{2}}\bar{\psi}\gamma^{\mu}\psi\bar{e}\gamma_{\mu}\gamma^{5}e$), the pseudo-tensor--tensor operator ($\mathcal{O}_{pt}\equiv$$\frac{1}{\Lambda^{2}}\bar{\psi}\sigma^{\mu\nu}i\gamma^{5}\psi\bar{e}\sigma_{\mu\nu}e$), the scalar--pseudo-scalar operator ($\mathcal{O}_{sp}\equiv\frac{1}{\Lambda^{2}}\bar{\psi}\psi\bar{e}i\gamma^{5}e$), and the pseudo-scalar--pseudo-scalar operator ($\mathcal{O}_{pp}\equiv\frac{1}{\Lambda^{2}}\bar{\psi}\gamma^{5}\psi\bar{e}\gamma^{5}e$) mediated interactions, where the constraints are loose $\sim$ $\mathcal{O}$(10) GeV. Again, the constraints apply to any DM having the considered effective interactions (Table 1).
@@
It is interesting to observe that collider searches have effectively ruled out a significant portion of the parameter space for DM effective interactions, while some leptophilic scenarios still remain viable and can be tested in the future FCC-ee. However, the exclusion of a region in the EFT framework in terms of $m_{\chi}$-$\Lambda$ does not imply that DM particles of mass $m_{\chi}$ are no longer viable. This exclusion essentially means that the interactions of the SM and DM via the heavy mediators (integrated out in the EFT prescription) are excluded. However, the viability of light mediators remains still valid, which may bring non-trivial changes in the results regarding the existence of lower mass DM. It definitely demands a dedicated study.